solve the following bondry value problems by the method of separation of variables

Ut = Uxx 0<x<1 , t>0

U(0,t)= 0 ,U(pi ,t)= 0

U(x , 0) = sinx + 2sin3x


1
Expert's answer
2020-12-28T11:57:55-0500

"\displaystyle
\frac{\partial u}{\partial t} = \frac{\partial^2 u}{\partial x^2}\\

\textsf{Let}\,\, u(x, t) = X(x)Y(t)\\

X''Y = XY' = 0,\,\, X''Y= XY'\\

X''Y= XY' = K\\

\frac{X''}{X} = \frac{Y'}{Y} = K\\

\frac{X''}{X} = K,\,\, \frac{Y'}{Y} = K\\


\textsf{If}\,\, K\,\, \textsf{is chosen to be negative}\\

\textsf{say}\,\, K = -\lambda^2,\,\,\textsf{then}\,\,\\

X'' = -\lambda^2 X\\

X = A\cos(\lambda x) + B\sin(\lambda x)\\

Y' = -\lambda^2 Y\\

Y = Ce^{-\lambda^2 t}\\

u(x,t) = \left( A\cos(\lambda x) + B\sin(\lambda x)\right)\left(Ce^{-\lambda^2 t}\right)\\

u(x,0) = C\left( A\cos(\lambda x) + B\sin(\lambda x)\right) = \sin(x) + 2\sin(3x)\\

u(0,0) = AC = 0\\

u(0,t) = ACe^{-\lambda^2 t} = 0\\

u(0,0) = AC = 0\\

\textsf{Avoiding trivial zeros, we select}\,\, A = 0 \\


u(\pi,t) = \left( A\cos(\pi\lambda) + B\sin(\pi\lambda)\right)\left(Ce^{-\lambda^2 t}\right) = 0\\

u(\pi,t) = \left(B\sin(\pi\lambda)\right)\left(Ce^{-\lambda^2 t}\right) = 0\\

\implies B\sin(\pi\lambda) = 0,\,\, \sin(\pi\lambda) = \sin(n \pi) \\

\therefore \lambda = n \\

u(x,t) = \left(B\sin(nx)\right)\left(Ce^{-n^2 t}\right) = D\sin(nx)e^{-n^2 t}\\

u(x,t)\,\,\textsf{can be superposed as;}\\

\begin{aligned}
u(x,t) &=\sum_{n = 1}^{\infty} D_n\sin(nx)e^{-n^2 t}
\end{aligned}\\

u(x,0) = \sin(x) + 2\sin(3x)\\

\begin{aligned}
u(x,0) = \sin(x) + 2\sin(3x) &=\sum_{n = 1}^{\infty} D_n\sin(nx)
\end{aligned}\\

\textsf{We have here a half-range Fourier sine}\\
\textsf{series representation of a function}\,\, u(x,0)\\
\textsf{defined over}\,\ 0< x < 1, \,\,\textsf{Extending}\,\, u(x,0)\,\, \textsf{as an}\\
\textsf{even periodic function with period}\,\, 2\,\, \textsf{and}\\
\textsf{using standard Fourier series theory gives}\\

\begin{aligned}
D_n &= 2\int_0^1 \left(\sin(x) + 2\sin(3x)\right) \sin(n x)\, \mathrm{d}x
\\&= \int_0^1 \left(\cos(x - nx) - \cos(x + nx) + 2\cos(3x - nx) - 2\cos(3x + nx)\right)
\\&= \left(\frac{\sin(x - nx)}{1 - n} - \frac{\sin(x + nx)}{1 + n} + \frac{2\sin(3x - nx)}{3 - n} - \frac{2\sin(3 + nx)}{3 + n}\right)\biggr\vert_0^1
\\&= \frac{\sin(1 - n)}{1 - n} - \frac{\sin(1 + n)}{1 + n} + \frac{2\sin(3 - n)}{3 - n} - \frac{2\sin(3 + n)}{3 + n}
\end{aligned}\\

\begin{aligned}
\therefore u(x,t) =\sum_{n = 1}^{\infty}\left(\frac{\sin(1 - n)}{1 - n} - \frac{\sin(1 + n)}{1 + n} + \frac{2\sin(3 - n)}{3 - n} - \frac{2\sin(3 + n)}{3 + n}\right)\sin(nx)e^{-n^2 t}\\
\end{aligned}\\"